OpenStudy (anonymous):

Carlos wants to deposit $900 into savings accounts at three different banks: Bank of Chance, Merchant Bank, and Utopian Financing. He will deposit two times as much into Merchant Bank as Bank of Chance because they offer a higher interest rate. He also expects the Utopian Financing deposit to be only 20% of the total of the other two deposits. How much will Carlos deposit into the Utopian Financing savings account? $150 $180 $250 $500

OpenStudy (arnavguddu):

let the deposit in Bank of Chance be X then deposit in Mar. Bank = 2X Total deposit till now = X+2X = 3X Thus, deposit in Uthopian Bank = 20 % of total = (20/100) 3X = 0.2 * 3X = 0.6X Total deposit = 3X + 0.6X = 3.6X = 900 So X = 900/3.6 Now deposit in Uthopian Bank = 0.6X = 0.6 * 900 / 3.6 = 6 * 900 / 36 = 900/6 = 150$